Factor Completely
a^3-4a^2+3a-12

Answer by sdmmadam@yahoo.com(530)

Factor Completely
a^3-4a^2+3a-12
=(a^3-4a^2)+(3a-12) by additive associativity
=a^2(a-4)+3(a-4)
=a^2p+3p where p=(a-4)
=p(a^2+3)
=(a-4)(a^2+3)
Answer: (a-4)(a^2+3)